# Classical emergence of the quantum-backreacted BTZ black hole from exponential electrodynamics

### v1

@@ I Introduction
In the absence of a fully developed theory of Quantum Gravity (QG), various strategies have been pursued to characterize the ultraviolet behavior of the gravitational field. For instance, quantum-backreacted spacetimes emerge when the effective energy-momentum tensor of a quantum field theory (QFT) is coupled to a classical set of field equations for the gravitational field. This semiclassical approach, though approximate, has proven valuable in capturing leading-order quantum effects in gravitational backgrounds, particularly around black holes.
@@ III Dynamical equivalence between quantum-corrected NMG and Einstein-NLED gravity
$\left.\left(T^{\text{NLED}}_{\mu\nu}-\frac{1}{2m^{2}}K_{\mu\nu}-\kappa N\langle T_{\mu\nu}\rangle\right)\right|_{g_{s}}=0.$ (27)
@@ IV Singularities and Horizon structure
In Fig. 1, we show examples of the behavior of the metric function $f$. We split the three distinct cases of critical points and where we can see the following characteristic forms: For the case of one critical point, the solution allocates only one horizon. For the case of two critical points given by the two real branches of the Lambert function, $f$ can manifest one, two or three horizons, depending on the relation between $c_{1}$ and $c_{3}$. Finally, for the degenerate single critical point, the solution has again only one horizon.
@@ VII Concluding remarks
Taken together, these results suggest that degenerate black hole geometries admitting both quantum-supported solutions and fully classical realizations can serve as concrete laboratories to explore the classical emergence of a semiclassical spacetime. In this setting, quantum corrections are effectively traded for nonlinearities in the matter sector, rather than for higher-derivative modifications of the gravitational dynamics, as is more commonly assumed. This viewpoint naturally points toward a broader correspondence between curvature corrections originating from quantum effects and classical nonlinear interactions, and raises the question of whether similar intersections of moduli spaces may occur beyond the static and purely electric sector. In this regard, it would be particularly interesting to search for quantum-supported duals of more general setups, such as nonlinearly charged rotating configurations or solutions accommodating a magnetic charge, like those reported in [30, 31, 32, 33], completely regular black holes as in [18, 19, 34, 35, 36, 37], or even related spacetimes in dimensions higher than three, for instance the analytical solutions reported in [38, 39, 40, 41, 42], to name a few representative examples.

### v2

@@ I Introduction
In the absence of a fully developed theory of Quantum Gravity (QG), various strategies have been pursued to characterize the ultraviolet behavior of the gravitational field. For instance, quantum-backreacted spacetimes emerge when the effective energy-momentum tensor of a quantum field theory (QFT) is coupled to a classical set of field equations for the gravitational field. This semiclassical approach, though approximate, has proven valuable in capturing leading-order quantum effects in gravitational backgrounds, particularly around black holes, including perturbative one-loop corrections to classical geometries and their phenomenological implications [1].
@@ III Dynamical equivalence between quantum-corrected NMG and Einstein-NLED gravity
$\left.\left(\kappa\,T^{\text{NLED}}_{\mu\nu}-\frac{1}{2m^{2}}K_{\mu\nu}-\kappa N\langle T_{\mu\nu}\rangle\right)\right|_{g_{s}}=0.$ (27)
@@ IV Singularities and Horizon structure
In Fig. 1, we display representative examples of the behavior of the metric function $f(r)$. We distinguish three qualitatively different cases according to the structure of its critical points. When $f$ admits a single critical point, the solution generically supports a single horizon. In the case of two critical points, arising from the two real branches of the Lambert function, the metric function may exhibit one, two, or three horizons, depending on the relative values of the parameters $c_{1}$ and $c_{3}$. Finally, for the degenerate configuration with a single (double) critical point, the solution again possesses a single horizon. It would be interesting to explore whether imposing suitable energy conditions on the energy-momentum tensor of the nonlinear electrodynamics further restricts the allowed configurations to the single-horizon case, in analogy with the mechanism recently identified in [26], where causal NLED theories with finite self-energy are shown to generically eliminate inner Cauchy horizons in charged black holes.
@@ VII Concluding remarks
Taken together, these results suggest that degenerate black hole geometries admitting both quantum-supported solutions and fully classical realizations can serve as concrete laboratories to explore the classical emergence of a semiclassical spacetime. In this setting, quantum corrections are effectively traded for nonlinearities in the matter sector, rather than for higher-derivative modifications of the gravitational dynamics, as is more commonly assumed. This view naturally points toward a broader correspondence between curvature corrections originating from quantum effects and classical nonlinear interactions, and raises the question of whether similar intersections of moduli spaces may occur beyond the static and purely electric sector. From a complementary perspective, it would also be interesting to investigate whether such classical mimics of semiclassical geometries admit a meaningful extension to black hole chemistry, where variations of the cosmological constant and additional parameters of quantum origin are reinterpreted holographically as changes in the number of degrees of freedom or in the holographic field theory scale, along the lines discussed in [32]. In this regard, it would be particularly interesting to search for quantum-supported duals of more general setups, such as nonlinearly charged rotating configurations or solutions accommodating a magnetic charge, like those reported in [33, 34, 35, 36], completely regular black holes as in [19, 20, 37, 38, 39, 40], or even related spacetimes in dimensions higher than three, for instance the analytical solutions reported in [41, 42, 43, 44, 45], to name a few representative examples.